# Local quantum detection of cosmological expansion: Unruh-DeWitt in spatially compact Milne

### v1

@@ IV.2.1 Oscillator modes
The response of the detector to the oscillatory modes shows similar asymptotic behaviour to periodic Minkowski, otherwise called the static cylinder universe. As the universe size approaches zero, the response of the detector to the oscillatory modes of the field (III.20) vanishes. This is behaviour that can be understood in terms of the energy of the oscillatory modes, since their minimum energy increases as the universe size decreases. As their minimum energy increases, the detector will be less able to exchange energy with these modes, and therefore its response to them will decrease. As the size of the universe approaches zero, their energy diverges and they invoke no response in the detector.
@@ VI Summary and discussion
Our detector was switched on and off sharply. Whilst this allows the duration of the interaction to be defined cleanly, it has the side effect that the response bears some imprint of the kicks at the switch-on and switch-off moments. In four or more spacetime dimensions, the effect from the kicks would in fact be divergent [40, 41, 42, 43]. A version of our analysis with a smoother switching could therefore be more readily comparable to that of a detector in four-dimensional spacetime. Another direction of work would be to compare the detector’s response with the field’s renormalised stress-energy tensor, and specifically the contributions to each from the zero mode, as done for the massive field in [6]. Finally, a simple way to introduce spacetime curvature would be to consider locally de Sitter spacetimes with Milne-like foliations, as done for the massive field in [8], providing insights into the special issues that arise with a massless minimally-coupled field in de Sitter spacetime [46, 21, 24]. We leave these questions subject to future work.
@@ Appendix B Large Detector Excitation Gap
It can be shown that $|f_{n}^{(1)}|\leq c_{1}\mu^{-1}|f_{n}^{(0)}|$ and $|f_{n}^{(2)}|\leq c_{2}\mu^{-1}|f_{n}^{(0)}|$ where $c_{1},c_{2}$ are independent of both $\mu$ and $n$, so for large positive $\mu$ we can take $|f_{n}|=|f_{n}^{(0)}|(1+\mathcal{O}(\mu^{-1}))$ uniformly in $n$.
@@
where the uniformity in $n$ allows us to extract the order bounding. This tells us that we can work with our leading order contribution $S^{(0)}$, but contributions from other terms will come into play at one order of $\mu$ lower. We can expand $S^{(0)}$ as
@@ B.1 Twisted Field
where $\psi$ and $\psi^{\prime}$ are the digamma [44, Eq 5.2.2] and trigamma [44, Eq 5.15.1] functions respectively, $\gamma$ is the Euler-Mascheroni constant arising from the known value of $\psi(1/2)$ [44, Eq 5.4.13], and the series expansion [44, Eq 5.7.6] has been used to obtain the trigamma. For large $\mu$, make use of expansions [47, Eq 2.4] and [44, Eq 5.15.8] to find
@@
$S_{2}^{(0)}(\mu)$ $=-\frac{1}{\pi\sigma_{f}}\mathrm{Re}\Biggl{[}\mathrm{e}^{i\mu\sigma_{f}}\mathrm{e}^{i\pi\lambda-\epsilon}\biggl{[}\alpha\mathcal{R}\left(\frac{\mu\xi\beta}{2\pi}+\frac{1}{2},\lambda+i\epsilon/\pi,1\right)$
@@
The Lipschitz-Lerch zeta functions in (B.12) can be expanded using [48, Eq. 7], which can be written as
@@ B.2 Untwisted Field
using the same expansions [47, Eq 1.3] and [44, Eq 5.15.8] for trigamma and digamma functions.
The untwisted expression for $S_{2}^{(0)}(\mu)$ continues in a similar fashion to the twisted field, by the introduction of a small positive imaginary component to $\lambda$ we find
$S_{2}^{(0)}(\mu)$ $=-\frac{\mu^{2}\xi^{2}\alpha\beta}{\pi}\mathrm{Re}\left[\sum_{n=1}^{\infty}\frac{\mathrm{e}^{i\mu\sigma_{f}}\left(\mathrm{e}^{2\pi i\lambda-\epsilon}\right)^{2}}{n(2\pi n+\mu\xi\alpha)(2\pi n+\mu\xi\beta)}\right]$
@@ Appendix C Large De-Excitation Gap
and since $\alpha<r_{n}<\beta$ for $S_{2}$, this integral contains a single minimum in the phase at $x=r_{n}$. Thus, methods of stationary phase such as [49, Eq. 3.2] can be used to find

### v2

@@ IV.1 Large Energy Gap
Note that the error term in (IV.2) should be understood as $\mathcal{O}\bigl{(}\omega^{-2}\ln|\omega/\omega_{0}|\bigr{)}$, where $\omega_{0}$ is an arbitrary positive constant of dimension inverse length, making the expression dimensionally consistent. The value of $\omega_{0}$ is of no consequence, since changing $\omega_{0}$ adds to $\omega^{-2}\ln|\omega/\omega_{0}|$ a multiple of $\omega^{-2}$, which is subdominant. We use a similar shorthand in error terms involving logarithms in the rest of the paper.
@@ IV.2.1 Oscillator modes
The response of the detector to the oscillatory field modes shows similar asymptotic behaviour to periodic Minkowski, otherwise called the static cylinder universe. As the universe size approaches zero, the response of the detector to the oscillatory modes of the field (III.20) vanishes. This is behaviour that can be understood in terms of the energy of the oscillatory modes, since their minimum energy increases as the universe size decreases. As their minimum energy increases, the detector will be less able to exchange energy with these modes, and therefore its response to them will decrease. As the size of the universe approaches zero, their energy diverges and they invoke no response in the detector.
@@ VI Summary and discussion
Our Unruh-DeWitt detector was switched on and off sharply. Whilst this allows the duration of the interaction to be defined cleanly, it has the side effect that the response bears some imprint of the kicks at the switch-on and switch-off moments. In four or more spacetime dimensions, the effect from the kicks would in fact be divergent [40, 41, 42, 43]. A version of our analysis with a smoother switching would therefore be more readily comparable to that of a detector in four-dimensional spacetime. A version of our analysis for a four-dimensional Milne universe with compact spatial sections would however be qualitatively different in that the universe size parameters would be discrete rather than continuous, because of the Mostow-Prasad rigidity theorem [46]. Further, a four-dimensional Milne universe has limited applicability in the description of our own universe, because of its constant expansion rate [11, 12, 13, 14, 15].
In four-dimensional cosmological spacetimes that model our own universe more closely, an Unruh-DeWitt detector as a late time probe of early universe phenomena has been considered in [7] in a setting with spatial flatness and $T^{3}$ spatial topology, with careful attention to the detector’s switching, in a communications channel setting in [47], and in a non-Markovian memory setting in [48]: all of these analyses could be generalised to explore the effect of the five continuous shape parameters that the $T^{3}$ spatial topology admits [49, 50]. All of these analyses could be further adapted to locally de Sitter cosmologies with $T^{3}$ spatial topology, as a model of the inflationary cosmological era, providing insights into the special issues that arise with a massless minimally-coupled field in de Sitter spacetime [51, 21, 24]. Another direction of work would be to compare the detector’s response with the field’s renormalised stress-energy tensor, and specifically the contributions to each from the zero mode, as done for the massive field in $1+1$ dimensions in [6]. We leave these questions subject to future work.
@@ Appendix B Large Detector Excitation Gap
It can be shown that $|f_{n}^{(1)}|\leq c_{1}\mu^{-1}|f_{n}^{(0)}|$ and $|f_{n}^{(2)}|\leq c_{2}\mu^{-1}|f_{n}^{(0)}|$ where $c_{1},c_{2}$ are independent of both $\mu$ and $n$. It follows that for large positive $\mu$ we have
$|f_{n}|=|f_{n}^{(0)}|(1+\mathcal{O}(\mu^{-1})),$ (B.4)
where the error term is independent of $n$.
@@
where the third equals sign follows from the $n$-independence of the error term in (B.4) and from the observation that $\left(1+\mathcal{O}(\mu^{-1})\right)^{2}=1+\mathcal{O}(\mu^{-1})$. This tells us that we can work with our leading order contribution $S^{(0)}$, but contributions from other terms will come into play at one order of $\mu$ lower. We can expand $S^{(0)}$ as
@@ B.1 Twisted Field
where $\psi$ and $\psi^{\prime}$ are the digamma [44, Eq 5.2.2] and trigamma [44, Eq 5.15.1] functions respectively, $\gamma$ is the Euler-Mascheroni constant arising from the known value of $\psi(1/2)$ [44, Eq 5.4.13], and the series expansion [44, Eq 5.7.6] has been used to obtain the trigamma. For large $\mu$, make use of expansions [52, Eq 2.4] and [44, Eq 5.15.8] to find
@@
$S_{2}^{(0)}(\mu)$ $=-\mathrm{Re}\Biggl{[}\frac{\mathrm{e}^{i\mu\sigma_{f}}\mathrm{e}^{i\pi\lambda-\epsilon}}{\pi\sigma_{f}}\biggl{[}\alpha\mathcal{R}\left(\frac{\mu\xi\beta}{2\pi}+\frac{1}{2},\lambda+i\frac{\epsilon}{\pi},1\right)$
@@
The Lipschitz-Lerch zeta functions in (B.13) can be expanded using [53, Eq. 7], which can be written as
@@ B.2 Untwisted Field
using the same expansions [52, Eq 1.3] and [44, Eq 5.15.8] for trigamma and digamma functions.
The untwisted expression for $S_{2}^{(0)}(\mu)$ continues similarly to the twisted field, by the introduction of a small positive imaginary component to $\lambda$ we find
$S_{2}^{(0)}(\mu)$ $=-\frac{\mu^{2}\xi^{2}}{\pi}\mathrm{Re}\left[\sum_{n=1}^{\infty}\frac{\alpha\beta\mathrm{e}^{i\mu\sigma_{f}}\left(\mathrm{e}^{2\pi i\lambda-\epsilon}\right)^{2}}{n(2\pi n+\mu\xi\alpha)(2\pi n+\mu\xi\beta)}\right]$
@@ Appendix C Large De-Excitation Gap
and since $\alpha<r_{n}<\beta$ for $S_{2}$, this integral contains a single minimum in the phase at $x=r_{n}$. Thus, methods of stationary phase such as [54, Eq. 3.2] can be used to find